OpenStudy (anonymous):

Whats the regression equation??? Height: 115, 124, 75, 75, 130, 74, 98, 87 Interval after min: 84, 78, 65, 69, 81, 72, 65, 66 Current eruption has a height of 92ft Use the significance level of 0.05. ŷ = ( ) + ( )x

OpenStudy (amistre64):

looks like a prime candidate for the ti83

OpenStudy (amistre64):

but then im not sure what the last line means for this

OpenStudy (amistre64):

input your values into list 1 and list 2 there is a function on your ti83 that will then give you the regression line.

OpenStudy (amistre64):

otherwise, it gets complicated, but excel or some spreadsheet makes life simpler for us

OpenStudy (anonymous):

ok I am trying now

OpenStudy (anonymous):

I cant get it

OpenStudy (anonymous):

This is my last question that I am having issues with

OpenStudy (amistre64):

yeah, its a bit complicated, but listing them is the key .... by hand, we would need to know more complicated results sum of x^2 sum of y^2 sum of xy sum of x sum of y average x average y and number of ordered pairs.

OpenStudy (anonymous):

I tried but I am getting the wrong data input

OpenStudy (anonymous):

Please please help me on this one

OpenStudy (amistre64):

ive got a spread sheet worked out hold on

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (amistre64):

ok, this is what im getting, and ill explain it

OpenStudy (anonymous):

Okay thank you bc that looks confusing lol

OpenStudy (amistre64):

we want to utilize the point slope formula, our point (avgX, avgY) and a slope that is determined by the formula: \[m=\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x}\] as such: \[\hat y=avgY+m(x-avgX)\] \[\hat y=avgY-m.avgX~+mx\] i used the spreadsheet to make all these calculations m = 2.356 avgY-m.avgX = 268.040

OpenStudy (amistre64):

now i dont know what your program is wanting as far as how many decimal places to rnd to

OpenStudy (amistre64):

i cut it off at 3 decimals

OpenStudy (amistre64):

and I have no idea what this is refering to: Use the significance level of 0.05

OpenStudy (anonymous):

So, it should then say ŷ = (2.356 ) + ( 268.04)x

OpenStudy (amistre64):

that is what i would use for it, but again, i dont know what your course uses as a rnding of decimals. I used 3 places.

OpenStudy (anonymous):

it says two decimal places

OpenStudy (amistre64):

ŷ = (avgY-m.avgX ) + (m)x ŷ = (268.04 ) + ( 2.36)x

OpenStudy (anonymous):

Ugh it said it wasn't right

OpenStudy (amistre64):

the equation itself is all i can answer to, the rest of it ... i havent a clue about.

OpenStudy (anonymous):

did you find the best predicted value?

OpenStudy (amistre64):

y^ is the best predictor

OpenStudy (amistre64):

when we formulate the regression line, we then use it for x=96

OpenStudy (amistre64):

well, 92 in this case

OpenStudy (amistre64):

if i had excel, it would allow me to verify the line, but i have something else on this computer ... guess i can use a website

OpenStudy (anonymous):

ŷ = (47.87 ) + ( 0.25)x what about from this?

OpenStudy (amistre64):

ok, an online calculator says ive got an error in the y intercept slope is fine tho

OpenStudy (amistre64):

got it now, i erred in multiplyng :) y = m(x-X) + Y y = mx -mX + Y ^^^^ i ignored the negative of it

OpenStudy (amistre64):

this is the correction y = 2.36x - 73.54